OpenStudy (anonymous):

find the lim as x approaches infinity of ((1) + (1/x))^kx where k is a real number. Hint: lim as x approaches infinity of ((1) + (1/x))^x = e I am leaning towards e^k but that's just a guess. How should I go about starting this problem?

OpenStudy (zarkon):

\[\left(1+\frac{1}{x}\right)^{kx}=\left[\left(1+\frac{1}{x}\right)^{x}\right]^k\]

OpenStudy (anonymous):

so that equals e^k?

OpenStudy (zarkon):

\[\lim_{x\to\infty}\left(1+\frac{1}{x}\right)^{kx}=\lim_{x\to\infty}\left[\left(1+\frac{1}{x}\right)^{x}\right]^k=\left[\lim_{x\to\infty}\left(1+\frac{1}{x}\right)^{x}\right]^k=e^k\]

OpenStudy (anonymous):

thank you so much Zarkon... would we apply the same use of exponents for something like this: lim from x to infinity of (1 + k/x)^xkprime where k is a real number?

OpenStudy (zarkon):

\[\lim_{x\to\infty}\left(1+\frac{k}{x}\right)^x\] let \(u=\frac{x}{k}\) then \(\frac{1}{u}=\frac{k}{x}\) and \(ku=x\) Assume k>0 \[\lim_{x\to\infty}\left(1+\frac{k}{x}\right)^x=\lim_{u\to\infty}\left(1+\frac{1}{u}\right)^{ku}=e^{k}\] similarly for k<0 if k=0 then we get \(1=e^{0}\)